# https://zielon.github.io/gem/

### v1

@@ Abstract
We present personalized Gaussian Eigen Models (GEMs) for human heads, a novel method that compresses dynamic 3D Gaussians into low-dimensional linear spaces. Our approach is inspired by the seminal work of Blanz and Vetter, where a mesh-based 3D morphable model (3DMM) is constructed from registered meshes. Based on dynamic 3D Gaussians, we create a lower-dimensional representation of primitives that applies to most 3DGS head avatars. Specifically, we propose a universal method to distill the appearance of a mesh-controlled UNet Gaussian avatar using an ensemble of linear eigenbasis. We replace heavy CNN-based architectures with a single linear layer improving speed and enabling a range of real-time downstream applications. To create a particular facial expression, one simply needs to perform a dot product between the eigen coefficients and the distilled basis. This efficient method removes the requirement for an input mesh during testing, enhancing simplicity and speed in expression generation. This process is highly efficient and supports real-time rendering on everyday devices, leveraging the effectiveness of standard Gaussian Splatting. In addition, we demonstrate how the GEM can be controlled using a ResNet-based regression architecture. We show and compare self-reenactment and cross-person reenactment to state-of-the-art 3D avatar methods, demonstrating higher quality and better control. A real-time demo showcases the applicability of the GEM representation.
Keywords:
Eigen Head Avatars, Gaussian Splatting
Figure 1: Given a multi-view video of a subject and meshes in full correspondence, we construct a high-quality Gaussian Eigen Model (GEM). Our method, employing a UNet (left), regresses a sequence of 3D Gaussians that undergo distillation to obtain compressed linear eigenbasis which we call Eigen Gaussians (right). To generate new photo-realistic 3D faces only a linear combination of the Gaussian basis has to be computed. We show that the required coefficients for the linear combination can be regressed from single images, allowing for real-time facial animation. The simplicity of a GEM leads to very efficient storage and rendering times.
@@ 1 Introduction
Half a century ago, Frederick Parke described a representation and animation technique to generate ,,animated sequences of a human face changing expressions” [48]. Using polygonal meshes, single facial expression states were described that could be combined with linear interpolation to generate new expression states (the ,,simplest way consistent with natural motion” [48]). Based on this principle, Blanz, and Vetter [3] introduced the so-called 3D morphable model (3DMM) - a statistical model of the 3D shape and appearance of human faces. Principle Component Analysis (PCA) is performed on a set of around 200 subjects that have been laser scanned and registered to a consistent template to find the displacement vectors (principal components) of how faces change the most, in terms of geometry and albedo. With this PCA basis, new faces can be generated by specifying the coefficients for the principle components taking a dot product of the coefficients with the basis to obtain offests, and adding them to the mean. State-of-the-art reports on face reconstruction and tracking [86], as well as on morphable models [8], state that this representation is widely used for performance capturing (regression-based and optimization-based) and builds the backbone of recent controllable photo-realistic 3D avatars that are equipped with neural rendering [63, 64, 81, 17, 11].
@@
Inspired by the simplicity of such mesh-based linear morphable models and addressing the lack of appearance realism of current 3DMMs, we propose a personalized linear appearance model based on 3D Gaussians as geometry primitive following 3D Gaussian Splatting (3DGS) [28]. In contrast to the concurrent work on Dynamic 3D Gaussian Avatars [38, 53, 83, 47, 79, 74, 59], our goal is a compact representation that does not need vast amounts of compute resources to generate novel expressions of the human. Unfortunately, most of the methods show that to produce high-quality results, one needs to employ heavy CNN-based architectures which are not well suited for commodity devices and tend to slow down the rendering pipeline. Moreover, those models comprise dozens of millions of parameters creating heavy checkpoints that can easily exceed 500 MB. This ultimately creates a major issue for distributing and managing personalized models. We tackle this problem by distilling a novel CNN-based architecture, leading to a personalized Gaussian Eigen Model, GEM in short. Our approach builds on Gaussian maps predicted from a modified UNet architecture [69] which is used for the UV space normalization required to properly build linear eigenbasis. Based on the per-subject trained model, we bootstrap the personalized GEM by computing an ensemble of linear basis on the predicted Gaussian maps of the training frames. Next, the basis is refined on the training corpus using photometric losses while preserving the orthonormality.
@@
This lightweight appearance basis is controlled with a relatively low number of parameters ranging from twenty up to fifty coefficients which can be specified w.r.t. the available compute resources and can for example be regressed by a ResNet-based model [9]. We demonstrate this for the scenarios of self-reenactment as well as cross-person animation. We showcase a real-time application of our method in the supplemental video.
In summary, our contributions are:
@@
    Gaussian Eigen Model (GEM) a universal distillation technique of 3D Gaussian head avatar model that is built upon an ensemble of eigenbasis.
@@
    A state-of-the-art CNN-based avatar appearance model that can be constructed using multi-view captures of a subject and serves as an example that can be used in combination with GEM.
  3. 3.
@@ 2 Related Work
The majority of current face representations and tracking techniques are based on parametric 3D morphable models (3DMM) [3, 37]. For a detailed overview of 3DMM-based reconstruction and appearance models, we refer to the state-of-the-art reports on face tracking and reconstruction [86], the report on morphable models [8], and the two neural rendering state-of-the-art reports [63, 64] that demonstrate how neural rendering can be leveraged for photo-realisitic facial or full body avatars. In the following, we review the recent methods that are focused on photo-realistic 3D avatars which build appearance models using neural radiance fields (NeRF) [43] or volumetric primitives like 3D Gaussians [28].
@@ 2.1 NeRF-based avatars
One of the first methods that combined 3DMM and NeRF is NeRFace [11]. In this work, a neural radiance field is directly conditioned by expression codes using the Basel Face Model (BFM) [3] optimized by the Face2Face tracker [67]. This idea gave rise to many methods [84, 81, 17, 82, 78, 76, 52, 12] following a similar approach, but attaching the radiance fields more explicitly to the surface of the 3DMM, e.g., by using the 3DMM-defined deformation field. To enable photorealistic results some methods employ StyleGAN2 [27] with a NeRF-based renderer [26, 1, 5]. Generative methods like EG3D [5] and PanoHead [1] employ a GAN-based training to predict triplane features that span a neural radiance field. GANAvatar [26] applies this scheme to reconstruct a personalized avatar.
Close to our method is StyleAvatar [69]. Based on 3DMM tracking the method learns a personalized avatar that benefits from a StyleUNet which incorporates StyleGAN [27] to decode the final image. Despite real-time capabilities, the method suffers from artifacts produced by the image-to-image translation network that we explicitly avoid by using Gaussian maps which can compensate for tracking misalignments by predicting corrective fields for the 3D Gaussians.
@@
Neural Volumes (NV) [39] started an impressive line of research in the context of digital avatars. Using multiview images with a Variational Auto-Encoder [30] and volumetric integration it encodes arbitrary dynamic scenes into a volume which can be deformed by traversing the latent code z. To better control the 3D space, Lombardi et al. [40] introduced Mixture of Volumetric Primitives (MVP) a hybrid representation based on primitives attached to the tracked mesh which ultimately replaced the encoder from NV. Each primitive is a volume represented as a small voxel with $32^{3}$ cells that on the vertices store RGB and opacity values. The final color is obtained by integrating values along a pixel ray. This hybrid representation became very popular spawning many follow-up projects [4, 55, 36, 62, 71]. The recently presented 3D Gaussians Splatting (3DGS) [28] introduced an alternative representation to MVP primitives. 3DGS represents a volume as a set of anisotropic 3D Gaussians, which are equivalently described as ellipsoids, in contrast to isotropic spheres used in Pulsar [35]. Similarly to DSS [70] it defines a backward pass of the rasterization technique Surface Splatting introduced by Zwicker et al. [87]. However, with the difference of using $\alpha$-blending [43] to integrate 3D Gaussians along a given ray. Numerous recent methods [83, 38, 59, 74, 80, 53, 10, 73, 24, 47, 32, 15] capitalize on the speed and quality of 3DGS. Qian et al. attached 3D Gaussians to the FLAME [37] mesh surface and applied a deformation gradient similar to Zielonka et al. [84] to orient the Gaussians according to the local Frenet frames of the surface. This method, however, does not utilize any information about expressions and, thus, struggles with pose-dependent changes (e.g., wrinkles, self-shadows) and, despite high-quality results, retrieves only a global static appearance model. Li et al. [38] are using the StyleUNet architecture [69] to regress front and back Gaussian maps. Employing the powerful CNN network on position maps, they achieve impressive results for human bodies. However, their architecture uses three separate StyleUNets each with two decoders for front and back texture which creates significant computational overhead making the whole pipeline slow. In contrast, we propose a more efficient CNN-based model that operates in a UV-space which we distill a lightweight linear 3D Gaussian Eigen Model (GEM).
@@ 3 Method
Current dynamic 3D Gaussian Avatars show unprecedented quality, however, they require sophisticated and often compute-heavy CNN-based architectures [38, 47, 74] leveraging large receptive fields to capture high-frequency details like pose-dependent wrinkles or self-shadows. We present an efficient CNN-based high-quality model and propose a technique to compress and distill such an avatar into an ensemble of eigenbasis that preserves its neural counterpart’s quality and expressiveness. The following section will explain our novel CNN-based appearance model and how to distill it into an alternative representation called the Gaussian Eigen Model (GEM).
Figure 2: CNN-based Gaussian Avatar Pipeline. Our CNN model produces delta Gaussian maps $\boldsymbol{D}_{i}$ [38, 47] and static color $\boldsymbol{C}$ from a multi-view video. Similarly to Animatable Gaussians [38], we constrain the network to operate in a reduced linear space, i.e., the mesh $\boldsymbol{M}_{i}$ is projected on a PCA basis $\boldsymbol{M}_{PCA_{i}}=\Gamma_{\text{PCA}}(\boldsymbol{M}_{i})$ which is then input to the network $\Pi$ after converting the mesh to a normal map $\phi(\boldsymbol{M}_{PCA_{i}})$.
Note that compared to other neural avatars [84, 11, 17, 39, 74, 69], GEM does not require a 3DMM [37, 50], i.e., training GEM only requires meshes in correspondence, while testing uses the GEM model parameters, making it effectively mesh and neural network free.
3.1 3D Gaussian Splatting Preliminaries
3.2 Appearance Maps Generator
As an example of a powerful Gaussian generator, we present a CNN-based appearance model inspired by Pix2Pix [22], which translates an input image to a specific output domain. In our case, Gaussian maps are generated from coarse normal maps of a tracked mesh. The idea of organizing the 3D Gaussians in 2D maps is used in several methods  [38, 47, 74]; each texel in such a map represents a 3D Gaussian with its parameters: offset vector for rotation, scale, position, and opacity. Similarly to StyleAvatar [69], we use a StyleGAN-based [27] encoder and decoder network for this image translation. However, in contrast to StyleAvatar, we propose a more lightweight image translation pipeline, where we reduce the number of encoders from three to two and the number of decoders from six to two, and decrease the size of the StyleGAN [27] decoder. To efficiently use the 2D map space, we do not use projective textures from meshes [69], but a UV parametrization which reduces half of the decoders in comparison to StyleAvatar. Moreover, we do not use triplets of the encoder-decoder, as we combine all properties of the Gaussians into one map $\boldsymbol{G}_{i}$ following ASH [47]. Therefore, we define our network as follows:
The final Gaussian map $\boldsymbol{G}_{i}\in\mathbb{R}^{11\times H\times W}$ is obtained by applying the deltas to the canonical Gaussians $\boldsymbol{G}$ [38, 83]. The deformed position of 3D means is computed as $\boldsymbol{T}_{i}(\boldsymbol{M}+\boldsymbol{D}_{i_{0:3}})$ (Section 3.3). Note that our color map C is static and does not model view-dependent effects; this means that we force the network to recover globally consistent colors for each Gaussian similar to a texture in the classic 3DMM. Therefore, $\boldsymbol{G}_{i}$ must model the wrinkles and self-shadows.
3.3 Deformation Model
The output of the CNN regressors $\Pi$ and $\Psi$ are in a canonical space. For the transformation from the canonical space to the deformed space, we employ deformation gradients, following Sumner et al. [60]. This approach allows for greater flexibility regarding input meshes, provided they maintain full correspondence. Given a mesh $\boldsymbol{M}_{PCA_{i}}$ for the frame $i$, we define the deformation gradients as $\mathbf{J}_{j}=\hat{\mathbf{E}}_{j}\mathbf{E}_{j}^{-1}$, where $\hat{\mathbf{E}}_{j}\in\mathbb{R}^{3\times 3}$ and $\mathbf{E}_{j}\in\mathbb{R}^{3\times 3}$ contain the Frenet frame (tangent, bi-tangent, normal) of the triangle $j$ defined in deformed and canonical spaces, respectively. Using these deformation gradients and the known correspondences between the Gaussian map and the meshes, we transform the Gaussians from the canonical space to the deformed space.
3.4 Implementation Details
where $\omega=0.2$, $\zeta=0.0075$ (after 150k iterations steps and zero otherwise), $\mathcal{L}_{\text{D-SSIM}}$ is a structural dissimilarity loss, and $\mathcal{L}_{\text{VGG}}$ is the perceptual VGG loss. $\mathcal{L}_{Reg}$ regularizes position offsets $\Delta_{pos_{j}}$ and scales $s_{scale_{j}}$ to stay small w.r.t. the input mesh. We train our model for $10^{6}$ steps using Adam [29] with a learning rate $5\mathrm{e}{-4}$ and a batch size of one which takes around 24h on a Nvidia A100.
Our method and all the baselines were trained using the same multiview input data sourced from the dataset provided by Qian [53], which includes multiview images from the NeRSamble dataset [33] as well as tracked meshes. These datasets are licensed and managed by MPI and TuD.
4 Gaussian Eigen Model (GEM)
Once a powerful appearance generator is available, we can build our universal eigenbasis model, GEM. For each training frame, we save the predicted Gaussian maps $\boldsymbol{G}_{i}$, where the positions are normalized by un-posing the head joint. The Gaussian maps could be successfully replaced with Gaussian point clouds, as long as their quantity remains unchanged. We stack the Gaussian maps in a matrix $\mathbf{A}\in\mathbb{R}^{N\times 11\times H\times W}$ where $N$ is the number of frames. Since the color $\boldsymbol{C}$ was optimized globally, it acts like a classic texture and therefore does not need distillation.

### v2

# Gaussian Eigen Models for Human Heads

@@ Abstract
Current personalized neural head avatars face a trade-off: lightweight models lack detail and realism, while high-quality, animatable avatars require significant computational resources, making them unsuitable for commodity devices. To address this gap, we introduce Gaussian Eigen Models (GEM), which provide high-quality, lightweight, and easily controllable head avatars. GEM utilizes 3D Gaussian primitives for representing the appearance combined with Gaussian splatting for rendering. Building on the success of mesh-based 3D morphable face models (3DMM), we define GEM as an ensemble of linear eigenbases for representing the head appearance of a specific subject. In particular, we construct linear bases to represent the position, scale, rotation, and opacity of the 3D Gaussians. This allows us to efficiently generate Gaussian primitives of a specific head shape by a linear combination of the basis vectors, only requiring a low-dimensional parameter vector that contains the respective coefficients. We propose to construct these linear bases (GEM) by distilling high-quality compute-intense CNN-based Gaussian avatar models that can generate expression-dependent appearance changes like wrinkles. These high-quality models are trained on multi-view videos of a subject and are distilled using a series of principle component analyses.
Once we have obtained the bases that represent the animatable appearance space of a specific human, we learn a regressor that takes a single RGB image as input and predicts the low-dimensional parameter vector that corresponds to the shown facial expression. We demonstrate that this regressor can be trained such that it effectively supports self- and cross-person reenactment from monocular videos without requiring prior mesh-based tracking. In a series of experiments, we compare GEM’s self-reenactment and cross-person reenactment results to state-of-the-art 3D avatar methods, demonstrating GEM’s higher visual quality and better generalization to new expressions. As our distilled linear model is highly efficient in generating novel animation states, we also show a real-time demo of GEMs driven by monocular webcam videos. The code and model will be released for research purposes.
Figure 1: We propose a method that represents 3D Gaussian head avatars in a network-free form as ensembles of eigenbases (GEM). Only a linear combination of these bases is needed to generate new primitives, which can be splatted using 3D Gaussian Splatting. We demonstrate that the necessary coefficients for a specific expression can be regressed from single images, enabling real-time facial animation and cross-reenactment. The simplicity of GEM results in highly efficient storage and rendering times.
@@ 1 Introduction
Half a century ago, Frederick Parke described a representation and animation technique to generate ,,animated sequences of a human face changing expressions” [53]. Using polygonal meshes, single facial expression states were described that could be combined with linear interpolation to generate new expression states (the ,,simplest way consistent with natural motion” [53]). Based on this principle, Blanz, and Vetter [3] introduced the so-called 3D morphable model (3DMM) - a statistical model of the 3D shape and appearance of human faces. Principle Component Analysis (PCA) is performed on a set of around 200 subjects that have been laser-scanned and registered to a consistent template to find the displacement vectors (principal components) of how faces change the most, in terms of geometry and albedo. With this PCA basis, new faces can be generated by specifying the coefficients for the principle components taking a dot product of the coefficients with the basis to obtain offsets, and adding them to the mean. State-of-the-art reports on face reconstruction and tracking [92], as well as on morphable models [8], state that this representation is widely used for facial performance capturing (regression-based and optimization-based) and builds the backbone of recent controllable photo-realistic 3D avatars that are equipped with neural rendering [67, 68, 86, 19, 12].
@@
Inspired by the simplicity of such mesh-based linear morphable models and addressing the lack of appearance realism of current 3DMMs, we propose a personalized linear appearance model based on 3D Gaussians as geometry primitive following 3D Gaussian Splatting (3DGS) [29]. In contrast to the work on Dynamic 3D Gaussian Avatars [41, 57, 90, 51, 84, 81, 63], our goal is a compact and light representation that does not need vast amounts of compute resources to generate novel expressions of the human. Unfortunately, most of the methods show that to produce high-quality results, one needs to employ heavy CNN-based architectures which are not well suited for commodity devices and tend to slow down the rendering pipeline. Moreover, those models comprise dozens of millions of parameters creating heavy checkpoints that can easily exceed 500 MB. This ultimately creates a major issue for distributing and managing personalized models. We tackle this problem by distilling a CNN-based architecture, leading to a personalized Gaussian Eigen Models for Human Heads, GEM in short. Our approach builds on Gaussian maps predicted from a modified UNet architecture [74] which is used for the UV space normalization required to properly build linear eigenbases. Based on the per-subject trained CNN model, we bootstrap the personalized GEM by computing an ensemble of linear bases on the predicted Gaussian maps of the training frames. The bases are refined on the training corpus using photometric losses while preserving their orthogonality.
@@
These lightweight appearance bases are controlled with a relatively low number of parameters ranging from twenty up to fifty coefficients which can be specified w.r.t. the available compute resources and can for example be regressed by a ResNet-based model [10]. We demonstrate this for the scenarios of self-reenactment as well as cross-person animation. We showcase a real-time application of our method in the supplemental video.
In summary, our main contributions are:
@@
    Gaussian Eigen Models for Human Heads (GEM) is a distillation technique of 3D Gaussian head avatar models built upon an ensemble of eigenbases.
@@ 2 Related Work
Figure 2: Given a multi-view video of a subject and mesh tracking, we create a dataset of 3D Gaussian point clouds for each frame in the sequence. Using this data, we distill a high-quality Gaussian Eigen Models for Human Heads (GEM). GEM is an ensemble of linear bases for each Gaussian primitive modality: position, opacity, scale, and rotation. Based on these bases, facial appearances are generated by a linear combination.
The majority of face representation and tracking techniques are based on parametric 3D morphable models (3DMM) [3, 40]. For a detailed overview, we refer to the state-of-the-art reports on face tracking and reconstruction [92], the report on morphable models [8], and the two neural rendering state-of-the-art reports [67, 68] that demonstrate how neural rendering can be leveraged for photo-realistic facial or full body avatars. Next, we review the recent methods for photo-realistic 3D avatars generation which build appearance models using neural radiance fields (NeRF) [46] or volumetric primitives like 3D Gaussians [29].
@@ 2.1 NeRF-based avatars
One of the first methods that combines a 3DMM and NeRF is NeRFace [12], where a neural radiance field is directly conditioned by expression codes of the Basel Face Model (BFM) [3, 71]. This idea gave rise to many methods [88, 86, 19, 87, 83, 80, 56, 13] following a similar approach, but attaching the radiance fields more explicitly to the surface of the 3DMM, e.g., by using the 3DMM-defined deformation field. For photorealistic results, some methods employ StyleGAN2-like architectures [28] with a NeRF-based renderer [27, 1, 5]. Generative methods like EG3D [5] and PanoHead [1] employ GAN-based training to predict triplane features that span a NeRF. GANAvatar [27] applies this scheme to reconstruct a personalized avatar.
Close to our method is StyleAvatar [74]. Based on 3DMM tracking the method learns a personalized avatar that benefits from a StyleUNet which incorporates StyleGAN [28] to decode the final image. Despite real-time capabilities, StyleAvatar suffers from artifacts produced by the image-to-image translation network that we explicitly avoid by using Gaussian maps which can compensate for tracking misalignments by predicting corrective fields for the 3D Gaussians.
@@
Using multiview images with a variational auto-encoder [31] and volumetric integration, Neural Volumes (NV) [42] encodes dynamic scenes into a volume which can be deformed by traversing a latent code z. To better control the 3D space, Lombardi et al. [43] introduce Mixture of Volumetric Primitives (MVP) a hybrid representation based on primitives attached to a tracked mesh which ultimately replaced the encoder from NV. Each primitive is a volume represented as a small voxel with $32^{3}$ cells that store RGB and opacity values. The final color is obtained by integrating values along a pixel ray. This hybrid representation inspired many follow-up projects [4, 59, 39, 66, 76]. As an alternative to MVP primitives, 3D Gaussian Splatting (3DGS) [29] represents a volume as a set of anisotropic 3D Gaussians, which are equivalently described as ellipsoids, in contrast to isotropic spheres used in Pulsar [36].
Numerous methods [90, 41, 63, 81, 85, 57, 11, 78, 25, 51, 34, 16, 91] capitalize on the speed and quality of 3DGS. Qian et al. [57] attach 3D Gaussians to the FLAME [40] mesh surface and apply a deformation gradient similar to Zielonka et al. [88] to orient the Gaussians according to the local Frenet frames of the surface. This method, however, does not utilize any information about expressions and, thus, struggles with pose-dependent changes (e.g., wrinkles, self-shadows) and, despite high-quality results, retrieves only a global static appearance model. Li et al. [41] use a StyleUNet-like CNN architecture [74] to regress front and back Gaussian maps. Employing the powerful CNN network on position maps, they achieve impressive results for human bodies with effects like pose-dependent wrinkle formation.
Please note that in this work, we focus on methods that directly output Gaussian primitives. This is an important distinction from a branch of methods that follow Deferred Neural Rendering [72], where a refinement CNN translates splatted features or coarse colors into the final image; for instance, Gaussian Head Avatars [81] and NGPA [16]. This distinction is important because Gaussian primitives cannot be fully distilled into an eigenbasis in this context, as the refinement CNN network is required to complete the rendering directly in the image space.
2.3 3DGS Compression Methods
Recently, several methods [52, 37, 38, 47, 9, 17] have been proposed to reduce the memory footprint of 3D Gaussian Splatting (3DGS). Papantonakis et al. [52] apply codebook quantization to the Gaussian primitive properties, alongside pruning of Spherical Harmonic (SH) coefficients based on their final contribution. In contrast to postprocessing approaches [52, 9, 38], Compact3D [47] employs a single-stage process that jointly optimizes both the codebook entries and the primitives. Fan et al. [9] calculate a significance score for each primitive by measuring its pixel hit count, thereby improving the pruning strategy. Most of these methods target static scenes or time-conditioned environments, unlike our approach, which focuses on efficient, fully controllable head avatars. Nonetheless, these compression techniques could be adapted to our animatable avatars to reduce memory usage.
@@ 3 Method
Recent dynamic 3D Gaussian Avatar methods show unprecedented quality, however, they require sophisticated and often compute-heavy CNN-based architectures [41, 51, 81] to capture high-frequency and dynamic details like pose-dependent wrinkles or self-shadows. The aim of this paper is to build on top of this quality but remove the compute-intense architecture during inference. Specifically, we propose to distill high-quality avatar models into lightweight linear animation models which we call GEMs. A GEM is defined by an ensemble of eigenbases that span the space of the 3D Gaussian primitives. These eigenbases are constructed via PCA applied on a dataset of per-frame Gaussian primitives, see Section 3.1.
An important distinction compared to other neural avatars [88, 12, 19, 42, 81, 74] is that GEM does not require a 3DMM [40, 55] at test time. We demonstrate that a GEM can be directly driven by a monocular video using a generalized image-based regression network, see Section 3.2.
3.1 Gaussian Eigen Model (GEM)
For our distillation, we reconstruct a sequence of normalized Gaussian primitives $\boldsymbol{D}=\{\boldsymbol{G}_{0},...,\boldsymbol{G}_{N-1}\}$. As input, we assume a multi-view video of the subject with $N$ time frames. Per time frame $i$, we reconstruct the 3D Gaussian pointcloud $\boldsymbol{G}_{i}$, where $\boldsymbol{G}_{i}$ contains the parameters that define the 3D Gaussians such as rotation $\text{rotation }\vec{\theta},\text{position }\vec{\phi},\text{opacity }\vec{\alpha},\text{scale }\vec{\sigma}$, and color $\text{color }\vec{c}$ such that $\boldsymbol{G}_{i}=\{\vec{\theta},\vec{\phi},\vec{\alpha},\vec{\sigma},\vec{c}\}$.
Reconstructing High-quality 3D Gaussian Primitives:
We are following the idea of organizing the 3D Gaussians in 2D maps [41, 51, 81, 63], where each pixel represents a 3D Gaussian with its parameters. We propose an adapted CNN-architecture of Animatable Gaussians (AG) [41], by merging the separate Style-U-Nets, reducing the convolutional layers, and operating in the UV space of the FLAME head model. In addition, we are employing deformation gradients following Sumner et al. [64] to handle the transformation from canonical to deformed space and treat the color as a global parameter. We refer to the suppl. mat. for a detailed explanation of the architectural changes. In comparison to the original AG model, our proposed CNN model produces slightly better results while being more efficient in terms of computing and memory. Using this model, we generate the per-frame Gaussian primitives $\boldsymbol{G}_{i}$ in the canonical space for all training time-frames. Note that for this reconstruction, we follow Animatble Gaussians and, thus, FLAME-based tracking is required. However, during inference, our model is independent of FLAME.
Distillation:
Given $\boldsymbol{D}=\{\boldsymbol{G}_{0},...,\boldsymbol{G}_{N-1}\}$, we build a personalized eigenbasis model, which is called GEM. We compute a statistical model for each Gaussian modality separately. Specifically, we create individual bases for rotation $\text{rotation }\boldsymbol{B}_{\theta},\text{position }\boldsymbol{B}_{\phi},\text{opacity }\boldsymbol{B}_{\alpha}$, and scale $\text{scale }\boldsymbol{B}_{\sigma}$ with respective means $\vec{\mu}_{\theta}$, $\vec{\mu}_{\phi}$, $\vec{\mu}_{\alpha}$ and $\vec{\mu}_{\sigma}$ via Principle Complonent Analysis (PCA) [26]. Note that the color $\boldsymbol{C}$ is optimized globally and, thus, acts as a classical texture without the need to apply PCA. A face model instance $\boldsymbol{G}$ is represented as a linear combination of these bases:
$\boldsymbol{G}=\left\{\vec{\mu}_{i}+\mathbf{B}_{i}\mathbf{k}_{i}\mid i\in\{\theta,\phi,\alpha,\sigma\},\vec{c}\right\},$ (1)
where $\mathbf{k}_{\theta}$, $\mathbf{k}_{\phi}$, $\mathbf{k}_{\theta}$ and $\mathbf{k}_{\sigma}\in\mathbb{R}^{M}$ are the linear coefficients which are defining the facial expression state, assuming $M$ principal components. As an example, Figure 3 shows position parameter $\mathbf{k}_{\phi}$ sampled in the range of $[-3\sigma_{\phi},3\sigma_{\phi}]$ ($\sigma_{\phi}$ being the std. deviation).
As the Gaussian primitives $\mathbf{D}$ might contain tracking failures and misalignments, the principle components $\boldsymbol{B}_{(\theta,\phi,\alpha,\sigma)}$ also contain artifacts as well. We, therefore, refine the bases using the training images directly, by applying a photometric reconstruction loss. We employ the same objectives from the CNN model training (see supp. mat).
$\mathcal{L}_{Color}=(1-\omega)\mathcal{L}_{1}+\omega\mathcal{L}_{\text{D-SSIM}}+\zeta\mathcal{L}_{\text{VGG}}$ (2)
We refine the base vectors for around 30k iterations. To ensure that the individual bases stay orthonormal throughout this refinement, every 1k steps, we orthogonalize the bases using QR decomposition. This refinement improves the training PSNR errors from 34.75dB to 36.68dB and 36.85dB for the training steps 0k, 5k, and 30k, respectively.